Answer what is asked in the given problem below. Show your solutions completely. Use an extra sheet of paper for your solutions. NO SOLUTIONS, NO POINTS.


The population is the weight of six pumpkins (in pounds) displayed in carnival "guess the weight" game booth. You are asked to guess the average weight of the six pumpkins by taking a random sample without replacement from the population. Suppose that the sample size of 5 pumpkins were drawn from this population (without replacement), describe the sampling distribution of the sample means.


Pumpkin


A


B


C


D


E


F


Weight in Pounds


19


14


15


9


10


17




What is the mean and variance of the sampling distribution of the sample means?


Compare these values to the mean and variance of the population.



1
Expert's answer
2022-06-10T11:08:50-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Pumpkin & Weight\ in\ pounds \\ \hline
 A & 19 \\
 \hdashline
 B & 14 \\
 \hdashline
 C & 15 \\
 \hdashline
 D & 9 \\
 \hdashline
 E & 10 \\
 \hdashline
 F & 17 \\
 \hdashline
\end{array}"

We have population values 9,10,14,15,17,19, population size N=6 and sample size n=5.

Mean of population "(\mu)" = "\dfrac{9+10+14+15+17+19}{6}=14"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{25+16+0+1+9+25}{6}""=\dfrac{38}{3}""\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{38}{3}}\approx3.56"

Select a random sample of size 5 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_5=6."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 9,10,14,15,17 & 65/5 \\
 \hdashline
 2 & 9,10,14,15,19 & 67/5 \\
 \hdashline
 3 & 9,10,14,17,19 & 69/5 \\
 \hdashline
 4 & 9,10,15,17,19 & 70/5\\
 \hdashline
 5 & 9,14,15,17,19 & 74/5 \\
 \hdashline
 6 & 10,14,15,17,19 & 75/4 \\
 \hdashline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 65/5 & 1/6 & 65/30 & 4225/150 \\
 \hdashline
 67/5 & 1/6 & 67/30 & 4489/150 \\
 \hdashline
 69/5 & 1/6 & 69/30 & 4761/150 \\
 \hdashline
 70/5 & 1/6 & 70/30 & 4900/150 \\
 \hdashline
 74/5 & 1/6 & 74/30 & 5476/150 \\
 \hdashline
 75/5 & 1/6 & 75/30 & 5625/150 \\
 \hdashline
\end{array}"


Mean of sampling distribution 



"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{420}{30}=14=\mu"


The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{29476}{150}-(14)^2=\dfrac{38}{75}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"